Determine the minimum angle of incidence for light traveling in flint glass in order to undergo total internal reflection from carbon disulfide. The indices of refraction respectively for flint glass and carbon disulfide are 1.66 and 1.63.
°
Answer #1

___________________________

The minimum angle of incidence for the light to undergo total internal reflection is critical angle.

n2 c = sin-1 ni

As the ray passes from flint glass (medium 1 with   n1 =1.66 ) to carbon disulfide (medium 2 with   n2 = 1.63)

n2 c = sin-1 ni  

= sin -1 (1.63/1.66) = 79.09 o

___________________________

Please rate

if any mistake in this answer please comment i will clarify your doubt . thank you